1. A ball is thrown horizontally from the top of a 60.0-m building and lands 100.0 m from the base of the building. Ignore air resistance. (a) How long is the ball in the air? (b) What must have been the initial horizontal component of the velocity? (c) What is the vertical component of the velocity just before the ball hits the ground? (d) What is the velocity (including both the horizontal and vertical components) of the ball just before it hits the ground? 


1
Expert's answer
2021-11-14T17:18:00-0500

(a) The time is


"t=\sqrt{2h/g}=3.5\text{ s}."

(b) The initial component:


"v_x=R/t=100/3.5=28.6\text{ m/s}."


(c) The vertical component of velocity:


"v_y=\sqrt{2gh}=34.3\text{ m/s}."

(d) The velocity on hitting the ground:


"v=\sqrt{v_x^2+v_y^2}=44.7\text{ m/s}."